10. Classify the behavior of f() = * at z = 20. If it is a zero or pole, give its order.
Answer #1

f(2)= Z 1+z² at z =0 () = 1/2 - It Yz² at z=0 YZ z²+1 Z2 = Z, a z=0 al z=0 z²+1 I al z= f(Yz) has zero - f (2) at z=oo has ze